help me pleaseeeeeeeeeeeeee!
lemon drops are to be mixed with jelly beans to make 100 pounds of mixture.
refer to the illustration and compute how many pounds of each candy should be used.
lemon drops $1.90 lb.
jelly beans $1.20 lb.
mixture$1.48 lb.

Answer by lwsshak3(11628)

lemon drops are to be mixed with jelly beans to make 100 pounds of mixture.
refer to the illustration and compute how many pounds of each candy should be used.
lemon drops $1.90 lb.
jelly beans $1.20 lb.
mixture$1.48 lb.
..
let x=lbs of lemon drops to use in mixture
then 100-x=lbs jelly beans to use in mixture
cost of lemon drops=1.9x
cost of jelly beans=1.2(100-x)
cost of the mixture=1.48*100
cost of lemon drops and jelly beans=cost of mixture
1.9x+1.2(100-x)=1.48*100
1.9x+120-1.2x=148
.7x=28
x=40lbs
100-x=100-40=60lbs
ans:
pounds of lemon drops to use in mixture=40
pounds of jelly beans to use in mixture=60